OpenStudy (anonymous):

The equation of line CD is y = -2x - 2. Write an equation of a line parallel to line CD in slope-intercept form that contains point (4, 5).

OpenStudy (anonymous):

y = -2x + 13 y = -1/2x + 7 y = -1/2x + 3 y = − 2x − 3

OpenStudy (anonymous):

the answer is y=-2x + 13